OpenStudy (anonymous):

(9x^2-4x-6)-(5x^2-4x+3)

OpenStudy (anonymous):

Ok, just subtract the like terms from the first parenthesis to the last. 9x^2 - 5x^2 = 4x^2 -4x - 4x = 8x -6 + 3 = -3 And then just put the terms together to get the final answer: 4x^2 + 8x - 3

OpenStudy (anonymous):

ok thanks that's what I got.

OpenStudy (dan815):

farmers are smart

OpenStudy (anonymous):

yw:)

OpenStudy (dan815):

can i buy your corn and wheat? how much

OpenStudy (anonymous):

Sorry. We only sell wholesale.

OpenStudy (dan815):

haha